let x represent one number and let y represent the other number. the sum of two numbers is 6. if the second number is subtracted from the other, the result is 0. find the numbers
Answer by stanbon(75887)

let x represent one number and let y represent the other number. the sum of two numbers is 6. if the second number is subtracted from the other, the result is 0. find the numbers
----
Equations:
x + y = 6
x - y = 0
---
2x = 6
x = 3
---
Since x+y = 6, y = 3
=========================
Cheers,
Stan H.
=========================